OpenStudy (anonymous):

Solve algebraically for x: (√x²+x-1)+11x = 7x+3?

OpenStudy (anonymous):

Well if that says the square root of x squared that is the same thing as x so replace the √x² with x. That leaves you with x+x-1+11x=7x+3 Then combine like terms (x, x, and 11x) to get 13x-1=7x+3 Then you need to get the x's on one side so subtract 7x from both sides to get 6x-1=3 Then you need to get everything else on the other side so add 1 to both sides to get 6x=4 Then to find x divide both sides by 6 to get x=4/6 Simplify that by dividing both the numerator and denominator by 2 and x=2/3

OpenStudy (arnavguddu):

note sqrt(x^2)=x.... so ur eq become... x+x-1 +11x=7x+3 or 13x-1=7x+3 change sides, bring all x to one side 13x-7x = 3+1 6x=4 x=4/6 x=2/3 done!!!